OpenStudy (anonymous):

108 is 36% of what number? Write and solve a proportion to solve the problem.

OpenStudy (31356):

If you consider that number is "X" and then solve x*36%=108 x=108/36% x=(108)/(36/100) x=300 So 300 is the number.